A 60 kg man wishes to run on ice. The coefficient of static friction between his shoes and the ice is .1. What is the maximum possible acceleration?
Answer #1

Maxmimum static friction = 0.1*60*9.8

Now the friction force is the force which will push him forwards else he will just slide.
So
F = 0.1*60*9.8
We know
F = ma
60*a = 0.1*60*9.8
a = 0.98 m/s^2